OpenStudy (anonymous):

(2qp^3)^3 Select one: a. 8q^3p^9 b. −8p^3q^9 c. m^3p^12p^6 d. 27q^3p^9

OpenStudy (mathstudent55):

The rule of raising a product to an exponent: Raise every factor to the exponent. Rule: \((abc)^n = a^nb^nc^n\) The rule of raising an exponent to an exponent: Multiply the exponents. Rule: \((a^m)^n = a^{m \times n} = a ^{mn}\) Example using both rules above: \((3x^2y^5z)^3 = 3^3 (x^2)^3(y^5)^3(z^1)^3 = 27x^{2 \times 3}y^{5 \times 3}z^{1 \times 3} = 27x^6y^{15}z^3\)